# Orbital Stability of Plane Waves in the Klein-Gordon Equation against Localized Perturbations

### v1

@@ 1.3. Discussion and outlook
Our spectral analysis in §7 shows that Theorem 4 is sharp in the sense that it holds up to the spectral stability boundary. That is, at the critical value $a^{2}f^{\prime}(a^{2})=2\max\{0,-f(a^{2})\}$ where condition (1.12) ceases to be satisfied, the plane wave (1.2) transitions from a spectrally stable to a spectrally unstable solution of (1.1). Notably, the condition (1.12) only involves the leading-order terms in the expansion (1.4) of the nonlinearity. When $f(a^{2})\geq 0$, the loss of stability corresponds to a transition from a defocusing to a focusing nonlinearity. In contrast, if $f(a^{2})<0$, the instability arises within the defocusing regime itself. We note that a change in stability is accompanied by a corresponding change in character (from elliptic to hyperbolic) of the Whitham’s modulation equation, which can be derived via a long-wave ansatz. For further details and the rigorous justification of the Whitham’s modulation equation for plane waves in the complex cubic Klein–Gordon equation, we refer to [HLS22].
Similar to the orbital stability result for time-periodic, spatially homogeneous solutions (1.7) to the NLS equation in [Zhi01], our result is locally uniform in space. Numerical simulations, see §6, suggest that an orbital stability result, where $H^{1}(x_{*}-R,x_{*}+R)$ in (1.17) is replaced by $H^{1}({\mathbb{R}})$ is not attainable. Specifically, our simulations reveal that localized perturbations trigger a local phase rotation of the plane wave, which subsequently spreads with constant speed in both spatial directions. That is, perturbations initiate a two-sided expanding phase front. Our numerical analysis indicates that the $L^{2}$-norm of $\vartheta(t)$ grows algebraically at rate $\smash{t^{\frac{1}{2}}}$, while its $L^{\infty}$-norm remains bounded. Thus, we do not expect the growth rate of the $L^{2}$-norm of $\vartheta(t)$ in Theorem 4 to be sharp. A compelling direction for future research would be to characterize the leading-order phase dynamics, which may, at least for long-wavelength perturbations, be governed by Whitham’s modulation equation; see [HLS22].
It remains an open question whether the conditions on $\theta_{\infty}$ in the regime $f(a^{2})\leq 0$ can be relaxed in Theorem 4, potentially admitting unbounded phase modulations in this case as well. Crossing the boundary $f(a^{2})=0$, however, marks a fundamental change in the character of equation (1.1), corresponding to a qualitative shift from a real to a tachyonic mass term in (1.4).
@@
$\begin{pmatrix}u_{n}(x,0)\\ \partial_{t}u_{n}(x,0)\end{pmatrix}=a\mathrm{e}^{\mathrm{i}kx-\mathrm{i}\theta_{\infty}(x)}\begin{pmatrix}1\\ \mathrm{i}\omega\end{pmatrix}+\begin{pmatrix}w_{n}(x)\\ v_{n}(x)\end{pmatrix},$
@@
We expect that the variational approach in this paper can be extended to higher spatial dimensions. In such settings, an intriguing direction is to explore whether the method can be combined with dispersive estimates to derive optimal or sharper bounds on the $\varrho$- and $\vartheta$-variables in Theorem 4. Owing to stronger dispersive decay in higher dimensions, one could even try to establish an asymptotic nonlinear stability result for plane-wave solutions. To the best of the authors’ knowledge, no asymptotic nonlinear stability results currently exist for periodic waves in dispersive Hamiltonian systems; we refer to [ARS24, Rod18] for linear asymptotic stability results for periodic solutions to the (generalized) Korteweg-de Vries equation. By contrast, the asymptotic nonlinear stability theory for soliton and kink solutions in such systems has seen significant advances over the past decades; see, for instance, the surveys [CM21, KMMn17] and references therein. Finally, another interesting avenue for future research is whether the orbital stability analysis presented here can be extended beyond plane waves to other classes of periodic solutions in Hamiltonian systems, particularly those that do not permit a reduction to a spatially homogeneous state.
@@ 2. Coordinate transformation and energy
Then, $\rho$ and $v$ measure the deviation from the manifold of equilibria =$\mathcal{M}=\{(\mathrm{e}^{\mathrm{i}\gamma},0)^{\top}:\gamma\in{\mathbb{R}}\}$ in (2.1), and $\phi$ tracks the solution’s orbit along $\mathcal{M}$ induced by gauge symmetry. Expressing
@@
it follows that the energy $E_{c}(w,v)$ is well-defined as long as $v\in L^{2}({\mathbb{R}})$ and $w=\mathrm{e}^{\rho+\mathrm{i}\phi}$ with $\rho\in H^{1}({\mathbb{R}})$ and $\phi^{\prime}\in L^{2}({\mathbb{R}})$. In case $c\omega+k\neq 0$, we must additionally require that $\phi^{\prime}$ is integrable, see Remark 6. In §4, we prove that this energy is indeed conserved and use (1.12) and (2.4) to derive lower and upper bounds which are sufficient to close an orbital stability argument.
Remark 6.
In the regime $f(a^{2})>0$, we can choose $c=-\frac{\omega}{k}\in(-1,1)$ so that the term
$\int_{\mathbb{R}}(c\omega+k)\phi^{\prime}(y)\mathrm{d}y$
in (2.4) vanishes. However, when $f(a^{2})\leq 0$, it turns out that this choice of $c$ no longer yields a good lower bound on $E_{c}(w,v)$, since (1.3) then implies $1-c^{2}=1-(\omega/k)^{2}\leq 0$. In this case, a different choice of $c$ is required, and we impose the additional condition in Theorem 4 that $\theta_{\infty}^{\prime}\in L^{1}({\mathbb{R}})$ is sufficiently small.
@@ 3. Local existence analysis
$u(x,t)=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t-\mathrm{i}\theta_{\infty}(x-ct)}$
to $w(y,t)=\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}$. We measure the deviation from the modulated plane wave by writing
$w(y,t)=\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}\left(1+z(y,t)\right).$ (3.1)
@@
$\qquad-2\mathrm{i}\theta_{\infty}^{\prime}\Big{(}\sqrt{1-c^{2}}\,\partial_{y}z_{1}+cz_{2}\Big{)}-(1+z_{1})\bigg{(}\sqrt{1-c^{2}}\big{(}\mathrm{i}\theta_{\infty}^{\prime\prime}+(\theta_{\infty}^{\prime})^{2}\bigg{)}-\frac{2(k+c\omega)}{\sqrt{1-c^{2}}}\theta_{\infty}^{\prime}\bigg{)}.$
@@
Note that $\Lambda$ is a skew-adjoint operator. As such, it generates a unitary group on the Hilbert space $X$ by Stone’s Theorem; see [EN00, Corollary II.3.24]. On the other hand, using that $H^{1}({\mathbb{R}})$ continuously embeds into $L^{\infty}({\mathbb{R}})$ and that we have $f\in{\mathcal{C}}^{1}({\mathbb{R}})$, one readily infers that $N$ is a well-defined locally Lipschitz continuous map. The result directly follows from standard semigroup theory for semilinear evolution equations; see [CH98, Theorem 4.3.4 and Proposition 4.3.9]. ∎
@@ 4. Energy estimates
In this section, we establish upper and lower bounds on the energy (2.2), which play a pivotal role in our stability argument. Expressing (2.2) in $Z$-coordinates via (3.1) and (3.2), we arrive at the quantity
[…]
where we recall that $U\in{\mathcal{C}}^{3}\big{(}[0,\infty),{\mathbb{R}}\big{)}$ is given by (2.3). Before showing that the energy is conserved and bounded from below, we first prove that it is bounded from above.
@@
Let $c\in(-1,1)$. There exists $C_{1}>0$ such that for each $\rho,\theta\in H^{1}({\mathbb{R}},{\mathbb{R}})$, $z_{1}\in H^{1}({\mathbb{R}},{\mathbb{C}})$, $z_{2}\in L^{2}({\mathbb{R}},{\mathbb{C}})$, and $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ satisfying $\theta_{\infty}^{\prime}\in L^{2}({\mathbb{R}})$, $(k+c\omega)\theta_{\infty}^{\prime}\in L^{1}({\mathbb{R}})$, and
@@
$\lvert E(z_{1},z_{2})\rvert\leq C_{1}\left(\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lvert k+c\omega\rvert\,\lVert\theta_{\infty}^{\prime}\rVert_{L^{1}({\mathbb{R}})}\right).$
@@
$\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}$ $\leq 2\lvert z_{1}^{\prime}(y)\rvert^{2}+2\lvert 1+z_{1}(y)\rvert^{2}\lvert\theta_{\infty}^{\prime}(y)\rvert^{2}$
@@
$\begin{split}\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)+\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}&=\big{(}\theta_{\infty}^{\prime}(y)-\theta^{\prime}(y)\big{)}\mathrm{e}^{2\rho(y)},\\ &=\theta^{\prime}(y)-\theta^{\prime}(y)\left(\mathrm{e}^{2\rho(y)}-1\right)+\theta_{\infty}^{\prime}(y)\mathrm{e}^{2\rho(y)}\end{split}$ (4.3)
@@
$|k+c\omega|\left\lvert\int_{\mathbb{R}}\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)+\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}\mathrm{d}y\right\rvert$
$\qquad=|k+c\omega|\left\lvert\int_{\mathbb{R}}\theta^{\prime}(y)\left(\mathrm{e}^{2\rho(y)}-1\right)-\theta_{\infty}^{\prime}(y)\mathrm{e}^{2\rho(y)}\mathrm{d}y\right\rvert$
$\qquad\leq C_{*}\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}\lVert\rho\rVert_{L^{2}({\mathbb{R}})}+|k+c\omega|\mathrm{e}^{2}\lVert\theta_{\infty}^{\prime}\rVert_{L^{1}({\mathbb{R}})},$
@@
Let $c\in(-1,1)$ and $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ be such that $\theta_{\infty}^{\prime}\in L^{2}({\mathbb{R}})$ and $(k+c\omega)\theta_{\infty}^{\prime}\in L^{1}({\mathbb{R}})$. Take $Z_{0}\in D(\Lambda)$. Let $Z\in{\mathcal{C}}([0,T),X)\cap{\mathcal{C}}^{1}([0,T),D(\Lambda))$ be the classical solution to (3.3) with initial condition $Z(0)=Z_{0}$, which was established in Proposition 7. Let $\tau\in(0,T]$ be such that the energy $E(Z(t))$ is well-defined for all $t\in[0,\tau)$. Then, we have
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\int_{\mathbb{R}}\left\lvert\partial_{y}\left((1+z(y,t))\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}\mathrm{d}y\right)$ $=\operatorname{Re}{\left\langle\partial_{y}\left((1+z(t)\mathrm{e}^{-\mathrm{i}\theta_{\infty}}\right),\partial_{y}\left(z_{t}(t)\mathrm{e}^{-\mathrm{i}\theta_{\infty}}\right)\right\rangle_{L^{2}({\mathbb{R}})}}$
$=-\operatorname{Re}{\left\langle\mathrm{e}^{\mathrm{i}\theta_{\infty}}\partial_{yy}\left((1+z(t))\mathrm{e}^{-\mathrm{i}\theta_{\infty}}\right),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{1}{2}\int_{\mathbb{R}}\theta_{\infty}^{\prime}(y)\lvert 1+z(y,t)\rvert^{2}\mathrm{d}y=\operatorname{Re}{\left\langle\theta_{\infty}^{\prime}(1+z(t)),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}E(Z(t))$ $=\operatorname{Re}{\left\langle 2cz_{ty}(t)-2\mathrm{i}(\omega+c\theta_{\infty}^{\prime})z_{t}(t),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}$
@@
Let $(a,k,\omega)\in(0,\infty)\times{\mathbb{R}}^{2}$ with $(k,\omega)\neq(0,0)$. Assume that the existence condition (1.3) and the spectral condition (1.12) hold. Then, there exist constants $C_{2}>0$, $\delta_{2}\in(0,1)$, and $c\in(-1,1)$ such that for each $\rho,\theta\in H^{1}({\mathbb{R}},{\mathbb{R}})$, $z_{1}\in H^{1}({\mathbb{R}},{\mathbb{C}})$, $z_{2}\in L^{2}({\mathbb{R}},{\mathbb{C}})$, and $a$-admissible $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ such that
@@
$\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\leq C_{2}\left(E(z_{1},z_{2})+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+|k+c\omega|\lVert\theta_{\infty}^{\prime}\rVert_{L^{1}({\mathbb{R}})}\right).$ (4.4)
@@
for $s\in[-1,1]$. Taking $\delta_{2}\in(0,1)$ such that $01-C_{*}\delta_{2}>$ 0, we use (4.5) and Young’s inequality to establish the pointwise lower bound
$\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}$ $=\mathrm{e}^{2\rho(y)}\left(\lvert\rho^{\prime}(y)\rvert^{2}+\lvert\theta^{\prime}(y)-\theta_{\infty}^{\prime}(y)\rvert^{2}\right)$
$\geq(1-C_{*}\delta_{2})\left(\lvert\rho^{\prime}(y)\rvert^{2}+(1-\delta_{2})\lvert\theta^{\prime}(y)\rvert^{2}-\frac{1-\delta_{2}}{\delta_{2}}\lvert\theta_{\infty}^{\prime}(y)\rvert^{2}\right)$
@@
$\begin{split}&\frac{(1-C_{*}\delta_{2})(1-\delta_{2})}{\delta_{2}}\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\int_{\mathbb{R}}\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{-\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}\mathrm{d}y\\ &\qquad\geq(1-C_{*}\delta_{2})\left(\lVert\rho^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+(1-\delta_{2})\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right).\end{split}$ (4.6)
@@
$\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)+\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}$
$\qquad=\theta^{\prime}(y)+2\theta^{\prime}(y)\rho(y)-\theta^{\prime}(y)\left(\mathrm{e}^{2\rho(y)}-1-2\rho(y)\right)+\theta_{\infty}^{\prime}(y)\mathrm{e}^{2\rho(y)}$
@@
$\begin{split}&\left\lvert\int_{\mathbb{R}}\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)+\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}\mathrm{d}y\right\rvert\\ &\qquad\leq\left(1+\frac{1}{2}C_{*}\delta_{2}\right)\left(\gamma\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}+\frac{1}{\gamma}\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}\right)+\mathrm{e}^{2}\|\theta_{\infty}^{\prime}\|_{L^{1}({\mathbb{R}})}.\end{split}$ (4.8)

### v2

@@ 1.3. Discussion and outlook
Our spectral analysis in §7 shows that Theorem 3 is sharp in the sense that it holds up to the spectral stability boundary. That is, at the critical value $a^{2}f^{\prime}(a^{2})=2\max\{0,-f(a^{2})\}$ where condition (1.12) ceases to be satisfied, the plane wave (1.2) transitions from a spectrally stable to a spectrally unstable solution of (1.1). Notably, the condition (1.12) only involves the leading-order terms in the expansion (1.4) of the nonlinearity. When $f(a^{2})\geq 0$, the loss of stability corresponds to a transition from a defocusing to a focusing nonlinearity. In contrast, if $f(a^{2})<0$, the instability arises within the defocusing regime itself. We note that a change in stability is accompanied by a corresponding change in character (from elliptic to hyperbolic) of the Whitham’s modulation equation, which can be derived via a long-wave ansatz. For further details and the rigorous justification of the Whitham’s modulation equation for plane waves in the complex cubic Klein–Gordon equation, we refer to [19].
Similar to the orbital stability result for time-periodic, spatially homogeneous solutions (1.7) to the NLS equation in [40], our result is locally uniform in space. Numerical simulations, see §6, suggest that an orbital stability result, where $H^{1}(x_{*}-R,x_{*}+R)$ in (1.16) is replaced by $H^{1}({\mathbb{R}})$ or where the right-hand side of (1.16) is replaced by $CE_{0}$, is not attainable. Specifically, our simulations reveal that localized perturbations trigger a local phase rotation of the plane wave, which subsequently spreads with constant speed in both spatial directions. That is, perturbations initiate a two-sided expanding phase front. Our numerical analysis indicates that the $L^{2}$-norm of $\vartheta(t)$ grows algebraically at rate $\smash{t^{\frac{1}{2}}}$, while its $L^{\infty}$-norm remains bounded. Thus, we do not expect the growth rate of the $L^{2}$-norm of $\vartheta(t)$ in Theorem 3 to be sharp. A compelling direction for future research would be to characterize the leading-order phase dynamics, which may, at least for long-wavelength perturbations, be governed by Whitham’s modulation equation; see [19].
@@
$\begin{pmatrix}u_{n}(x,0)\\ \partial_{t}u_{n}(x,0)\end{pmatrix}=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\theta_{\infty}(x)}\begin{pmatrix}1\\ \mathrm{i}\omega\end{pmatrix}+\begin{pmatrix}w_{n}(x)\\ v_{n}(x)\end{pmatrix}$,
@@
We expect that the variational approach in this paper can be extended to higher spatial dimensions. In such settings, an intriguing direction is to explore whether the method can be combined with dispersive estimates to derive optimal or sharper bounds on the $\varrho$- and $\vartheta$-variables in Theorem 3. Owing to stronger dispersive decay in higher dimensions, one could even try to establish an asymptotic nonlinear stability result for plane-wave solutions. To the best of the authors’ knowledge, no asymptotic nonlinear stability results currently exist for periodic waves in dispersive Hamiltonian systems; we refer to [1, 32] for linear asymptotic stability results for periodic solutions to the (generalized) Korteweg-de Vries equation. By contrast, the asymptotic nonlinear stability theory for soliton and kink solutions in such systems has seen significant advances over the past decades; see, for instance, the surveys [10, 26] and references therein. Finally, another interesting avenue for future research is whether the orbital stability analysis presented here can be extended beyond plane waves to other classes of periodic solutions in Hamiltonian systems, particularly those that do not permit a reduction to a spatially homogeneous state.
@@ 2. Coordinate transformation and energy
Then, $\rho$ and $v$ measure the deviation from the manifold of equilibria $\mathcal{M}=\{(\mathrm{e}^{\mathrm{i}\gamma},0)^{\top}:\gamma\in{\mathbb{R}}\}$ in (2.1), and $\phi$ tracks the solution’s orbit along $\mathcal{M}$ induced by gauge symmetry. Expressing
@@
it follows that the energy $E_{c}(w,v)$ is well-defined as long as $v\in L^{2}({\mathbb{R}})$ and $w=\mathrm{e}^{\rho+\mathrm{i}\phi}$ with $\rho,\phi\in H^{1}({\mathbb{R}})$. In §4, we slightly modify this energy to allow for unbounded initial phase modulations. Moreover, we prove that this modified energy is conserved and use (1.12) and (2.4) to derive lower and upper bounds which are sufficient to close an orbital stability argument.
@@ 3. Local existence analysis
$u(x,t)=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t+\mathrm{i}\theta_{\infty}(x-ct)}$
to $w(y,t)=\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}$. We measure the deviation from the modulated plane wave by writing
$w(y,t)=\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\left(1+z(y,t)\right).$ (3.1)
@@
$\qquad+2\mathrm{i}\theta_{\infty}^{\prime}\Big(\sqrt{1-c^{2}}\,\partial_{y}z_{1}+cz_{2}\Big)+(1+z_{1})\bigg(\sqrt{1-c^{2}}\big(\mathrm{i}\theta_{\infty}^{\prime\prime}-(\theta_{\infty}^{\prime})^{2}\bigg)-\frac{2(k+c\omega)}{\sqrt{1-c^{2}}}\theta_{\infty}^{\prime}\bigg).$
@@
Note that $\Lambda$ is a skew-adjoint operator. As such, it generates a unitary group on the Hilbert space $X$ by Stone’s Theorem; see [13, Corollary II.3.24]. On the other hand, using that $H^{1}({\mathbb{R}})$ continuously embeds into $L^{\infty}({\mathbb{R}})$ and that we have $f\in{\mathcal{C}}^{1}({\mathbb{R}})$, one readily infers that $N$ is a well-defined locally Lipschitz continuous map. The result directly follows from standard semigroup theory for semilinear evolution equations; see [7, Theorem 4.3.4 and Proposition 4.3.9]. ∎
@@ 4. Energy estimates
In this section, we slightly modify the energy (2.2) and establish upper and lower bounds, which play a pivotal role in our stability argument. Expressing (2.2) in $Z$-coordinates via (3.1) and (3.2) and correcting the integrand with $2(k+c\omega)\theta_{\infty}^{\prime}$, we arrive at the quantity
$\begin{split}E(z_{1},z_{2})&=\frac{1}{2}\int_{\mathbb{R}}(1-c^{2})\left(\lvert z_{2}(y)\rvert^{2}+\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}\right)+U\left(\lvert 1+z_{1}(y)\rvert^{2}\right)\\ &\qquad-\,2(k+c\omega)\left(\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)-\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}+\theta_{\infty}^{\prime}(y)\right)\mathrm{d}y,\end{split}$ (4.1)
where we recall that $U\in{\mathcal{C}}^{3}\big([0,\infty),{\mathbb{R}}\big)$ is given by (2.3). The correction term is chosen so as to eliminate linear terms in $\theta_{\infty}^{\prime}$ from the integrand, which need not be integrable under the sole assumption $\theta_{\infty}^{\prime}\in L^{2}({\mathbb{R}})$. Before showing that the energy is conserved and bounded from below, we first prove that it is indeed well-defined and bounded from above.
@@
Let $c\in(-1,1)$. There exists $C_{1}>0$ such that for each $\rho,\theta\in H^{1}({\mathbb{R}},{\mathbb{R}})$, $z_{1}\in H^{1}({\mathbb{R}},{\mathbb{C}})$, $z_{2}\in L^{2}({\mathbb{R}},{\mathbb{C}})$, and $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ satisfying $\theta_{\infty}^{\prime}\in L^{2}({\mathbb{R}})$ and
@@
$\lvert E(z_{1},z_{2})\rvert\leq C_{1}\left(\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right).$
@@
$\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}$ $\leq 2\lvert z_{1}^{\prime}(y)\rvert^{2}+2\lvert 1+z_{1}(y)\rvert^{2}\lvert\theta_{\infty}^{\prime}(y)\rvert^{2}$
@@
$\begin{split}&\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)-\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}+\theta_{\infty}^{\prime}(y)\\ &\qquad=\theta_{\infty}^{\prime}(y)-\big(\theta_{\infty}^{\prime}(y)+\theta^{\prime}(y)\big)\mathrm{e}^{2\rho(y)}=-\theta^{\prime}(y)-\left(\theta_{\infty}^{\prime}(y)+\theta^{\prime}(y)\right)\left(\mathrm{e}^{2\rho(y)}-1\right)\end{split}$ (4.3)
@@
$|k+c\omega|\left\lvert\int_{\mathbb{R}}\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)-\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}+\theta_{\infty}^{\prime}(y)\mathrm{d}y\right\rvert$
$\qquad=|k+c\omega|\left\lvert\int_{\mathbb{R}}\left(\theta_{\infty}^{\prime}(y)+\theta^{\prime}(y)\right)\left(\mathrm{e}^{2\rho(y)}-1\right)\mathrm{d}y\right\rvert$
$\qquad\leq C_{*}\left(\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}\right)\lVert\rho\rVert_{L^{2}({\mathbb{R}})},$
@@
Let $c\in(-1,1)$ and $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ be such that $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$. Take $Z_{0}\in D(\Lambda)$. Let $Z\in{\mathcal{C}}([0,T),X)\cap{\mathcal{C}}^{1}([0,T),D(\Lambda))$ be the classical solution to (3.3) with initial condition $Z(0)=Z_{0}$, which was established in Proposition 5. Let $\tau\in(0,T]$ be such that the energy $E(Z(t))$ is well-defined for all $t\in[0,\tau)$. Then, we have
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\int_{\mathbb{R}}\left\lvert\partial_{y}\left((1+z(y,t))\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}\mathrm{d}y\right)$ $=\operatorname{Re}{\left\langle\partial_{y}\left((1+z(t)\mathrm{e}^{\mathrm{i}\theta_{\infty}}\right),\partial_{y}\left(z_{t}(t)\mathrm{e}^{\mathrm{i}\theta_{\infty}}\right)\right\rangle_{L^{2}({\mathbb{R}})}}$
$=-\operatorname{Re}{\left\langle\mathrm{e}^{-\mathrm{i}\theta_{\infty}}\partial_{yy}\left((1+z(t))\mathrm{e}^{\mathrm{i}\theta_{\infty}}\right),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{1}{2}\int_{\mathbb{R}}\theta_{\infty}^{\prime}(y)\lvert 1+z(y,t)\rvert^{2}-\theta_{\infty}^{\prime}(y)\mathrm{d}y=\operatorname{Re}{\left\langle\theta_{\infty}^{\prime}(1+z(t)),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}E(Z(t))$ $=\operatorname{Re}{\left\langle 2cz_{ty}(t)-2\mathrm{i}(\omega-c\theta_{\infty}^{\prime})z_{t}(t),z_{t}(t)\right\rangle_{L^{2}({\mathbb{R}})}}$
@@
Let $(a,k,\omega)\in(0,\infty)\times{\mathbb{R}}^{2}$ with $(k,\omega)\neq(0,0)$. Assume that the existence condition (1.3) and the spectral condition (1.12) hold. Then, there exist constants $C_{2}>0$, $\delta_{2}\in(0,1)$, and $c\in(-1,1)$ such that for each $\rho,\theta\in H^{1}({\mathbb{R}},{\mathbb{R}})$, $z_{1}\in H^{1}({\mathbb{R}},{\mathbb{C}})$, $z_{2}\in L^{2}({\mathbb{R}},{\mathbb{C}})$, and $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ such that $\theta_{\infty}^{\prime}\in L^{2}({\mathbb{R}})$ and
@@
$\lVert z_{2}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\rho\rVert_{H^{1}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\leq C_{2}\left(E(z_{1},z_{2})+\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right),$ (4.4)
@@
for $s\in[-1,1]$. Taking $\delta_{2}\in(0,1)$ such that $1-C_{*}\delta_{2}>0$, we use (4.5) to establish the pointwise lower bound
$\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}$ $=\mathrm{e}^{2\rho(y)}\left(\lvert\rho^{\prime}(y)\rvert^{2}+\lvert\theta^{\prime}(y)+\theta_{\infty}^{\prime}(y)\rvert^{2}\right)$
$\geq(1-C_{*}\delta_{2})\left(\lvert\rho^{\prime}(y)\rvert^{2}+\lvert\theta^{\prime}(y)+\theta_{\infty}^{\prime}(y)\rvert^{2}\right)$
@@
$\begin{split}&\int_{\mathbb{R}}\left\lvert\partial_{y}\left((1+z_{1}(y))\mathrm{e}^{\mathrm{i}\theta_{\infty}(y)}\right)\right\rvert^{2}\mathrm{d}y\geq(1-C_{*}\delta_{2})\left(\lVert\rho^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}+\lVert\theta^{\prime}+\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right).\end{split}$ (4.6)
@@
$\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)-\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}+\theta_{\infty}^{\prime}(y)$
$\qquad=-\theta^{\prime}(y)-2\rho(y)\left(\theta^{\prime}(y)+\theta_{\infty}^{\prime}(y)\right)-\left(\mathrm{e}^{2\rho(y)}-1-2\rho(y)\right)\left(\theta_{\infty}^{\prime}(y)+\theta^{\prime}(y)\right)$
@@
$\begin{split}&\left\lvert\int_{\mathbb{R}}\operatorname{Im}\left((1+z_{1}(y))\overline{z_{1}^{\prime}(y)}\right)-\theta_{\infty}^{\prime}(y)\lvert 1+z_{1}(y)\rvert^{2}+\theta_{\infty}^{\prime}(y)\mathrm{d}y\right\rvert\\ &\qquad\leq\left(1+\frac{1}{2}C_{*}\delta_{2}\right)\left(\gamma\lVert\rho\rVert_{L^{2}({\mathbb{R}})}^{2}+\frac{1}{\gamma}\lVert\theta^{\prime}+\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}^{2}\right).\end{split}$ (4.8)